The following table shows the levels of retirement benefits given to a group of
workers in a given establishment.

Retirement benefits £ ‘000 No of retirees (f)
20 – 29 50
30 – 39 69
40 – 49 70

Required:-
Calculate the percentile coefficient of Kurtosis and briefly comment on the value
obtained
50 – 59 90
60 – 69 52
70 – 79 40
80 – 89 11
1
Expert's answer
2019-06-27T12:48:59-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Class \ Intervals & Frequencies\ (f_i) & Cumulative\ Frequencies \\ \hline
 20-29 & 50 & 50 \\
30-39 & 69 & 119 \\
40-49 & 70 & 189 \\
50-59 & 90 & 279 \\
60-69 & 52 & 331 \\
70-79 & 40 & 371 \\
80-89 & 11 & 382 \\
\end{array}""\sum f_i=382"

The percentile coefficient of Kurtosis, denoted as "K_p," is defined in terms of quartiles and percentiles as


"K_p={Q_3-Q_1 \over 2(P_{90}-P_{10})}"

"Q_1=L_1+\bigg[{\sum f_i\times 1/4-C_f \over f_1}\bigg]\cdot C_1"

"\sum f_i\times 1/4=382\times 1/4=95.5, L_1=30,""C_f=50, C_1=10, f_1=69"

"Q_1=30+\bigg[{95.5-50 \over 69}\bigg]\cdot 10=36.5942"

"Q_3=L_3+\bigg[{\sum f_i\times 3/4-C_f \over f_3}\bigg]\cdot C_3"

"\sum f_i\times 3/4=382\times 3/4=286.5, L_3=60,""C_f=279, C_3=10, f_3=52"

"Q_3=60+\bigg[{286.5-279 \over 52}\bigg]\cdot 10=61.4423"

"P_k=L_1+\bigg[{\sum f_i\times (k/100)-C_f \over f_k}\bigg]\cdot C_k"

"k=90, \sum f_i\times (k/100)=382\times (90/100)=343.8,""L_1=70, C_f=331, C_{90}=10, f_{90}=40"

"P_{90}=70+\bigg[{343.8-331 \over 40}\bigg]\cdot 10=73.20"

"k=10, \sum f_i\times (k/100)=382\times (10/100)=38.2,""L_1=20, C_f=0, C_{10}=10, f_{10}=50"


"P_{10}=10+\bigg[{38.2-0 \over 50}\bigg]\cdot 10=17.64"

The percentile coefficient of Kurtosis


"K_p={61.4423-36.5942 \over 2(73.20-17.64)}=0.2236"

Since "K_p=0.2236<0.263", the curve is leptokurtic or thin.